What is the diagonal of a rectangular room which measures 15 feet by 17 feet?

Answer by checkley79(3341)

A^2+B^2=C^2
15^2+17^2=C^2
225+289=C^2
514=C^2
C=SQRT514
C=22.67 ANS.